OpenStudy (anonymous):

A large moving box has a volume of 20 cubic meters. The width of the box is 0.5 meter. The length and the height of the box are each whole number measurements that are greater than 3 meters. What could be the possible dimensions of the box? The box has a length of 66 meters, height of meters, and width of 0.5 meter.

OpenStudy (anonymous):

@Awolflover1

OpenStudy (anonymous):

20 = 0.5 * l * h So then: 40 = l * h What two numbers can multiply to 40?

OpenStudy (anonymous):

2 and 20

OpenStudy (anonymous):

Try another set of numbers, since they have to be bigger than 3.

OpenStudy (anonymous):

4 and 10

OpenStudy (anonymous):

There you go. :)

OpenStudy (anonymous):

ok thank you